# Enhancing Large Language Models in Coding Through Multi-Perspective Self-Consistency

### v2

@@ Abstract
Large language models (LLMs) have exhibited remarkable ability in code generation. However, generating the correct solution in a single attempt still remains a challenge. Prior works utilize verification properties in software engineering to verify and re-rank solutions in a majority voting manner. But the assumption behind them that generated verification properties have better qualities than solutions may not always hold. In this paper, we treat them equally as different perspectives of LLMs’ reasoning processes. We propose the Multi-Perspective Self-Consistency (MPSC) framework incorporating both inter- and intra-consistency across outputs from multiple perspectives. Specifically, we prompt LLMs to generate diverse outputs from three perspectives, Solution, Specification and Test case, constructing a 3-partite graph. With two measure functions of consistency, we embed both inter- and intra-consistency information into the graph. The optimal choice of solutions is then determined based on analysis in the graph. MPSC significantly boosts performance of foundation models (ChatGPT in this paper) on various benchmarks, including HumanEval (+15.91%), MBPP (+6.43%) and CodeContests (+9.37%), even surpassing GPT-4.
@@ 1 Introduction
We evaluate MPSC on four widely used code generation benchmarks, including HumanEval (Chen et al., 2021), HumanEval+ (Liu et al., 2023), MBPP (Austin et al., 2021) and CodeContests (Li et al., 2022). Experimental results show that our method boosts the performance of ChatGPT by a large margin, 15.91% in HumanEval, 15.64% in HumanEval+, 6.43% in MBPP and 9.37% in CodeContests [^1] We will release our code on GitHub.. Our framework even surpasses GPT-4 (OpenAI, 2023) as shown in Figure 1.
@@ 2.4 Intra-Consistency within the Same Perspective
, where $C$ is the normalizing constant so that measures of outputs in one perspective sum up to $1$.
@@ 3.2 Main Results
Table 3: The performance of MPSC-Semantic with different foundation models. ${}^{\dagger}$: We use nucleus sampling with temperature as 0.2 instead of greedy generation in this experiment. The best performance is shown in bold.
@@ 4 Related Work
Based on Chain-of-thought mechanism (Wei et al., 2022), previous works have adopted various prompting techniques and decoding strategies to reveal the consistency of LLM outputs and further enhance the capabilities of LLMs. One line of approaches decodes multiple times from the same perspective and aggregate the results (Wang et al., 2022; Zhou et al., 2022; Jung et al., 2022; Sun et al., 2022). For example, Wang et al. (2022) targets tasks with fixed answer sets and scores each answer based on the output frequency. Building on this, Sun et al. (2022) introduces recitation as context for augmentation. While Jung et al. (2022) focus on the two-value entailment relations (True or False) between statements and explanations. They treat the inference process as a weighted MAX-SAT problem and utilize a logistic solver to solve it. Another line draws inspiration from the “Dual Process" in cognitive science, which posits that human reasoning is dominated by System 1 and System 2 (Daniel, 2017; Sloman, 1996). As a result, these approaches require LLMs to play different roles like generator (i.e. System 1) and verifier (i.e. System 2), and optimize the result iteratively by a conversational way (Madaan et al., 2023; Shinn et al., 2023; Zhu et al., 2023). Xiong et al. (2023a) also proposes the concept of "inter-consistency". Instead of referring to the consistency within the same LLM, they focus to tackle the inter-inconsistency problem between different models with a formal debate framework.
@@
LLMs pretrained on large-scale code data have demonstrated strong capabilities in the field of code generation, such as Codex (Chen et al., 2021), AlphaCode (Li et al., 2022), CodeGen (Nijkamp et al., 2023), InCoder (Fried et al., 2022), StarCoder (Li et al., 2023), Code Llama (Rozière et al., 2023), WizardCoder (Luo et al., 2023), DeepSeekCoder (Guo et al., 2024). However, they remain unreliable, particularly in scenarios involving complex input-output mappings. Because of the low tolerance of compilers and operating systems for bugs, the instability makes LLMs hard to deploy into real-world applications. Several methods have been proposed to mitigate the phenomenon (Shi et al., 2022; Chen et al., 2022; Zhang et al., 2022; Key et al., 2022; Ni et al., 2023; Dong et al., 2023; Olausson et al., 2023; Chen et al., 2023; Zhang et al., 2023). The line of work with the most direct relevance to ours is to re-rank generated solutions in a post-hoc manner. For example, Shi et al. (2022) matches the execution results of generated solutions for minimum Bayes risk selection. Zhang et al. (2022) prompts another model as a reviewer to check whether generated programs satisfy the given language instruction by measuring $p(\text{instruction}|\text{program})$. CodeT (Chen et al., 2022) additionally generates test cases to verify the generated solutions. Similarly, ALGO (Zhang et al., 2023) additionally generates exhaustive search algorithms as oracle programs to generate high quality test cases for verification.
@@ Appendix A Details of the Iterative Algorithm
       ${\bm{f}}^{(0)}\longleftarrow{\bm{y}}$ ${\bm{T}}\longleftarrow{\bm{D}}^{-\frac{1}{2}}{\bm{W}}{\bm{D}}^{-\frac{1}{2}}$ $i\longleftarrow 0$ Do
             ${\bm{f}}^{(i+1)}\longleftarrow\alpha{\bm{T}}{\bm{f}}^{(i)}+(1-\alpha){\bm{y}}$ $i\longleftarrow i+1$
      While $||{\bm{f}}^{(i)}-{\bm{f}}^{(i-1)}||\leq\epsilon$${\bm{f}}^{*}\longleftarrow{\bm{f}}^{(i)}$ return ${\bm{f}}^{*}$
@@ Appendix C Discussion about Pass@$k$
Appendix D More Analysis
D.1 MPSC with User-provided Test Cases
@@
D.2 Analysis of Other Perspectives
@@
D.3 Time Overhead
@@ Appendix F Prompt for MPSC
I want you to act as a python programmer. Given a docstring about a python method, you need to write its pre-conditions in one test function “def preconditions(input)" and post-conditions in another test function “def postconditions(input, output):". You should ensure invalid input or output of the method will raise error in the two test functions. ‘‘‘Python {Demonstration Docstrings 1} pass #Pre-conditions {Demonstration Preconditions 1} #Post-conditions {Demonstration Postconditions 1} ‘‘‘ ‘‘‘Python {Demonstration Docstrings 2} pass #Pre-conditions {Demonstration Preconditions 2} #Post-conditions {Demonstration Postconditions 2} ‘‘‘ ‘‘‘Python {Docstrings} pass
[h!] I want you to act like a Python programmer. I will give you the declaration of a function and comments about its property. You need to implement the body of the function in the code block. Do not modify any code I provide. Do not provide any explanations. Here is the question. ‘‘‘Python {Docstrings} ‘‘‘
‘‘‘Python # Given a docstring, continue to write the following code with 10 valid assertion statements to check the correctness of the function. Provide diverse test cases. {Docstrings} pass # check the correctness of with 10 different valid assertion statements in the form of “assert {entry point}(…) == …" assert

### v3

@@ Abstract
Large language models (LLMs) have exhibited remarkable ability in code generation. However, generating the correct solution in a single attempt still remains a challenge. Prior works utilize verification properties in software engineering to verify and re-rank solutions in a majority voting manner. But the assumption behind them that generated verification properties have better qualities than solutions may not always hold. In this paper, we treat them equally as different perspectives of LLMs’ reasoning processes. We propose the Multi-Perspective Self-Consistency (MPSC) framework incorporating both inter- and intra-consistency across outputs from multiple perspectives. Specifically, we prompt LLMs to generate diverse outputs from three perspectives, Solution, Specification and Test case, constructing a 3-partite graph. With two measure functions of consistency, we embed both inter- and intra-consistency information into the graph. The optimal choice of solutions is then determined based on analysis in the graph. MPSC significantly boosts performance of foundation models (ChatGPT in this paper) on various benchmarks, including HumanEval (+15.91%), MBPP (+6.43%) and CodeContests (+9.37%), even surpassing GPT-4.111  The code is available at https://github.com/skpig/MPSC.
@@ 1 Introduction
We evaluate MPSC on four widely used code generation benchmarks, including HumanEval (Chen et al., 2021), HumanEval+ (Liu et al., 2023), MBPP (Austin et al., 2021) and CodeContests (Li et al., 2022). Experimental results show that our method boosts the performance of ChatGPT by a large margin, 15.91% in HumanEval, 15.64% in HumanEval+, 6.43% in MBPP and 9.37% in CodeContests. Our framework even surpasses GPT-4 (OpenAI, 2023) as shown in Figure 1.
@@ 2.4 Intra-Consistency within the Same Perspective
, where $C$ is the normalizing constant so that measures of outputs in one perspective sum up to $1$, $K(v_{i})$ represents the other outputs within the same perspective as $v_{i}$.
@@ 3.2 Main Results
Table 3: The performance of MPSC-Semantic with different foundation models. : We use nucleus sampling with temperature as 0.2 instead of greedy generation in this experiment. The best performance is shown in bold.
@@ 4 Related Work
Based on Chain-of-thought mechanism (Wei et al., 2022), previous works have adopted various prompting techniques and decoding strategies to reveal the consistency of LLM outputs and further enhance the capabilities of LLMs. One line of approaches decodes multiple times from the same perspective and aggregate the results (Wang et al., 2022; Zhou et al., 2022; Jung et al., 2022; Sun et al., 2022; Chen et al., 2023a). For example, Wang et al. (2022) targets tasks with fixed answer sets and scores each answer based on the output frequency. Building on this, Sun et al. (2022) introduces recitation as context for augmentation. While Jung et al. (2022) focus on the two-value entailment relations (True or False) between statements and explanations. They treat the inference process as a weighted MAX-SAT problem and utilize a logistic solver to solve it. Another line draws inspiration from the “Dual Process" in cognitive science, which posits that human reasoning is dominated by System 1 and System 2 (Daniel, 2017; Sloman, 1996). As a result, these approaches require LLMs to play different roles like generator (i.e. System 1) and verifier (i.e. System 2), and optimize the result iteratively by a conversational way (Madaan et al., 2023; Shinn et al., 2023; Zhu et al., 2023). Xiong et al. (2023a) also proposes the concept of "inter-consistency". Instead of referring to the consistency within the same LLM, they focus to tackle the inter-inconsistency problem between different models with a formal debate framework.
@@
LLMs pretrained on large-scale code data have demonstrated strong capabilities in the field of code generation, such as Codex (Chen et al., 2021), AlphaCode (Li et al., 2022), CodeGen (Nijkamp et al., 2023), InCoder (Fried et al., 2022), StarCoder (Li et al., 2023), Code Llama (Rozière et al., 2023), WizardCoder (Luo et al., 2023), DeepSeekCoder (Guo et al., 2024). However, they remain unreliable, particularly in scenarios involving complex input-output mappings. Because of the low tolerance of compilers and operating systems for bugs, the instability makes LLMs hard to deploy into real-world applications. Several methods have been proposed to mitigate the phenomenon (Shi et al., 2022; Chen et al., 2022; Zhang et al., 2022; Key et al., 2022; Ni et al., 2023; Dong et al., 2023; Olausson et al., 2023; Chen et al., 2023b; Zhang et al., 2023). The line of work with the most direct relevance to ours is to re-rank generated solutions in a post-hoc manner. For example, Shi et al. (2022) matches the execution results of generated solutions for minimum Bayes risk selection. Zhang et al. (2022) prompts another model as a reviewer to check whether generated programs satisfy the given language instruction by measuring $p(\text{instruction}|\text{program})$. CodeT (Chen et al., 2022) additionally generates test cases to verify the generated solutions. Similarly, ALGO (Zhang et al., 2023) additionally generates exhaustive search algorithms as oracle programs to generate high quality test cases for verification.
@@ 5 Conclusion
Acknowledgments
This work was supported by National Key R&D Program of China (2021YFF0901502), Beijing Science and Technology Program (Z231100007423011) and Key Laboratory of Science, Technology and Standard in Press Industry (Key Laboratory of Intelligent Press Media Technology). We sincerely thank Xinyu Hu, Mingqi Gao, Xiao Pu and Xiang Chen for providing insightful advice about this work. We also appreciate the anonymous reviewers for their helpful comments.
@@ Appendix A Details of the Iterative Algorithm
       ${\bm{f}}^{(0)}\longleftarrow{\bm{y}}$
       ${\bm{T}}\longleftarrow{\bm{D}}^{-\frac{1}{2}}{\bm{W}}{\bm{D}}^{-\frac{1}{2}}$
       $i\longleftarrow 0$
       Do
             ${\bm{f}}^{(i+1)}\longleftarrow\alpha{\bm{T}}{\bm{f}}^{(i)}+(1-\alpha){\bm{y}}$
             $i\longleftarrow i+1$
      While $||{\bm{f}}^{(i)}-{\bm{f}}^{(i-1)}||\leq\epsilon$
      ${\bm{f}}^{*}\longleftarrow{\bm{f}}^{(i)}$
       return ${\bm{f}}^{*}$
@@ Appendix C Discussion about Pass@$k$
C.1 Analysis of Other Perspectives
@@
Benchmark HumanEval HumanEval+
Metric Pass@1 Pass@2 Pass@5 Pass@1 Pass@2 Pass@5
WizzardCoder-34B 67.84 72.12 75.98 58.70 62.88 66.88
+CodeT 73.17+5.33 73.17+1.05 72.03-3.95 64.21+5.51 64.36+1.48 63.41-3.47
Code Llama-34B 51.78 59.24 67.07 41.49 48.30 55.93
+CodeT 67.99+16.21 68.17+8.93 68.28+1.21 55.26+13.77 56.70+8.4 57.90+1.97
WizzardCoder-13B 60.35 66.10 72.01 50.25 56.00 61.98
+CodeT 66.86+6.51 67.18+1.08 67.93-4.08 58.23+7.98 58.72+2.72 58.99-2.99
Code Llama-13B 44.63 50.99 57.86 35.93 41.71 48.19
+CodeT 57.99+13.36 58.25+7.26 57.91+0.05 50.03+14.1 50.46+8.75 50.48+2.29
WizzardCoder-7B 53.81 59.62 66.06 45.06 50.83 57.69
+CodeT 63.17+9.36 63.36+3.74 63.41-2.65 54.13+9.07 55.05+4.22 55.74-1.95
Code Llama-7B 39.38 45.18 52.79 34.33 39.18 45.25
+CodeT 51.68+12.30 51.83+6.65 51.90-0.89 44.06+9.73 44.47+5.29 44.71-0.54
Table 12: The performance of MPSC-Semantic with different foundation models. : We use nucleus sampling with temperature as 0.2 instead of greedy generation in this experiment. The best performance is shown in bold.
Benchmark HumanEval HumanEval+ MBPP CodeContest
Metric Pass@$1$ Pass@$2$ Pass@$5$ Pass@$1$ Pass@$2$ Pass@$5$ Pass@$1$ Pass@$2$ Pass@$5$ Pass@$1$ Pass@$2$ Pass@$5$
GPT-3.5-Turbo 68.38 76.24 83.15 58.75 66.58 73.96 66.80 72.34 76.60 2.57 4.22 7.16
GPT-4 81.48 86.31 90.46 70.52 75.48 79.54 71.26 74.27 76.99 6.1 8.28 11.72
MPSC 81.5 83.7 91.05 72.18 75.2 81.22 72.78 74.24 78.25 10.77 12.64 13.84
Table 13: The performance of MPSC with 100 solutions, 50 specifications and 50 test cases
Appendix D More Analysis
D.1 MPSC with User-provided Test Cases
D.2 Time Overhead
@@ D.3 Stability of MPSC
D.4 MPSC with Limited API Calls
We here discuss another setting, where MPSC utilizes 100 solutions, 50 specifications and 50 test cases, requiring identical API calls to the foundation model baselines. The results shown in Table 13 again prove the supreme performance of MPSC over baselines under fair comparison.
D.5 Comparison on Other Foundation Models
Table 3 demonstrates the generalization of MPSC on other foundation models. We also conduct an experiment to compare MPSC with CodeT, the strongest baseline, under this setting. The results are presented in Table 12.
@@ Appendix F Prompt for MPSC
Prompt for Generating Specifications I want you to act as a python programmer. Given a docstring about a python method, you need to write its pre-conditions in one test function “def preconditions(input)" and post-conditions in another test function “def postconditions(input, output):". You should ensure invalid input or output of the method will raise error in the two test functions. ‘‘‘Python {Demonstration Docstrings 1} pass #Pre-conditions {Demonstration Preconditions 1} #Post-conditions {Demonstration Postconditions 1} ‘‘‘ ‘‘‘Python {Demonstration Docstrings 2} pass #Pre-conditions {Demonstration Preconditions 2} #Post-conditions {Demonstration Postconditions 2} ‘‘‘ ‘‘‘Python {Docstrings} pass
[h!] Prompt for Generating Solutions I want you to act like a Python programmer. I will give you the declaration of a function and comments about its property. You need to implement the body of the function in the code block. Do not modify any code I provide. Do not provide any explanations. Here is the question. ‘‘‘Python {Docstrings} ‘‘‘
Prompt for Generating Test Cases ‘‘‘Python # Given a docstring, continue to write the following code with 10 valid assertion statements to check the correctness of the function. Provide diverse test cases. {Docstrings} pass # check the correctness of with 10 different valid assertion statements in the form of “assert {entry point}(…) == …" assert